OpenStudy (anonymous):

If y= lnx prove that dy/dx=1/x

OpenStudy (alekos):

by definition.... int(1/x) = lnx

OpenStudy (unklerhaukus):

y = ln x e^y = x now differentiate both sides with respect to x

OpenStudy (ikram002p):

use the definition of f'(x)= lim h-->0 ([f(x+h)- f(h)/h)

OpenStudy (ikram002p):

so u have f(x)=ln x \( f'(x) = \lim_{h \rightarrow 0}\frac{f(x+h)-f(h)}{h}\)

OpenStudy (ikram002p):

\( f'(x)=\lim_{h \rightarrow 0}\frac{\ln (x+h)- \ln h} {h}\)

OpenStudy (unklerhaukus):

e^y = x e^y dy/dx = 1